A constant force of 25 N is applied as shown to a block which undergoes a displacement of 7.5 m to the right along a frictionless surface while the force acts. What iswork done by the force?

W=Fcos(30)s
W=25cos(30)7.5=160J

What else do I need to do?????
Answer #1

force F = 25 N, displacement s = 7.5 m
work = Fs cosθ = 162.37 J

answered by: confused by adverbs and adjectives
Answer #2

W=F.S

=25*7.5

=187.5 J

answered by: Lady Don
Answer #3

work is said to be done when force produces displacement
and if force is Fcosθ

and displacement is d

then work done = Fcosθ * d

answered by: Kurgi
Answer #4

Net work done = Work done by the force + Work done by the friction(=0)

Work done by force = F * x * cosθ

your question doesn't mention anyθ but your answer does so I'm pretty confused here.

-> for noθ, Workdone = 25 * 7.5 = 187.5 J

-> forθ= 30 (as per your answer), Workdone = 25 * 7.5 * cos30 = 162.37 (NOT 160)

answered by: Katy Marshall
Answer #5
your answer is correct.
answered by: janey b